# Twin branching in shape memory alloys: a 1D model with energy dissipation effects

### v3

@@ Abstract
We develop a 1D model of twin branching in shape memory alloys. The free energy of the branched microstructure comprises the interfacial and elastic strain energy contributions, both expressed in terms of the average twin spacing treated as a continuous function of the position. The total free energy is then minimized, and the corresponding Euler–Lagrange equation is solved numerically using the finite element method. The model can be considered as a continuous counterpart of the recent discrete model of Seiner et al. (2020), and our results show a very good agreement with that model in the entire range of physically relevant parameters. Furthermore, our continuous setting facilitates incorporation of energy dissipation into the model. The effect of rate-independent dissipation on the evolution of the branched microstructure is thus studied. The results show that significant effects on the microstructure and energy of the system are expected only for relatively small domain sizes.
@@ 1 Introduction
The theoretical analysis of twin branching started with the seminal contributions of Kohn and Müller (1992, 1994), followed by many related works, mostly in the mathematics community (e.g., Conti, 2000; Capella and Otto, 2009; Chan and Conti, 2015; Dondl et al., 2016; Simon, 2021). This line of research delivers, in particular, the scaling laws for energy and for characteristic dimensions of the branched microstructure. As a follow up, a quantitative model of twin branching in shape memory alloys has been recently developed by Seiner et al. (2020). The model assumes that the branched microstructure consists of a family of self-similar branching generations, as in several earlier models, and refined estimates of the elastic strain energy of branching have been derived for the considered microstructure. The parameters characterizing the actual microstructure are then obtained by minimizing the total free energy. This model will be used as a reference for our model developed in this work.
Direct spatially-resolved simulation of a branched microstructure by using, for example, the phase-field method, is not feasible because of the multiscale nature of the branched microstructure. However, some features characteristic for twin branching can be observed in phase-field simulations (e.g., Finel et al., 2010; Tůma and Stupkiewicz, 2016; Amini et al., 2023). These are limited to tip splitting of martensite needles, or the like, and thus correspond to just a single branching generation. To the best of our knowledge, spatially-resolved simulations of hierarchical branched microstructures have not been reported to date. A kind of hierarchical twin branching is reported in the related work of Li and Luskin (1999), but the interfacial energy is not included in that model (only the elastic strain energy is minimized) so that the model does not capture the essential competition between the interfacial and elastic strain energy contributions.
@@
We call our model a continuous model since the twin spacing is treated as a continuous function of the position and the governing equation can be formulated as an ODE. This is in contrast to the model of Seiner et al. (2020), which we call a discrete model, since that model considers a discrete number of branching generations thus leading to a step-like dependence of the twin spacing on the position. Moreover, the solution is obtained from a discrete minimization problem, in particular, involving minimization with respect to the (integer) number of branching generations.
A model similar to our model has been recently presented by Zhang et al. (2024) for the analysis of experimental data on austenite–martensite interfacial patterns in NiMnGa reported in the reference. That model also considers the twin spacing as a continuous function of the position and adopts the same form of the free energy density in the branched martensite. The difference is that our model includes the energy of elastic micro-strains at the A–MM interface, which is disregarded in the model of Zhang et al. (2024). This leads to different boundary conditions for the corresponding ODE. The model of Zhang et al. (2024), once fitted to their experimental data, correctly describes the profile of the twin spacing in the vicinity of the austenite–martensite interface.
@@
The remainder of the paper is organized as follows. In Section 2, the discrete model of twin branching of Seiner et al. (2020) is briefly outlined as a reference for our subsequent developments. In Section 3, the continuous model of twin branching is developed and sample results of finite-element computations are presented, including a comparison with the predictions of the model of Seiner et al. (2020). As a side note, an analytical solution to this (static) problem is discussed in A. In Section 4, energy dissipation is introduced into the model and the corresponding effects are studied for a few representative cases of evolution of the branched microstructure. Technical details concerning the derivation of the evolution problem and its computational treatment are presented in B and C, respectively. A general discussion on the model and its applicability is presented in Section 5, followed by concluding remarks in Section 6.
@@ 2.1 Austenite–twinned martensite interface
where $\hat{{\rm\bf a}}$ is a non-zero vector and ${\rm\bf R}$ is a rotation tensor. Consider further a simple laminate formed by the two martensite variants with $\lambda$ denoting the volume fraction of the variant ${\rm\bf U}_{A}$. The overall compatibility of the austenite and twin laminate has then the form of the habit plane equation,
@@ 2.2 Discrete model of twin branching
Assuming that $L_{i}\gg h_{i}$, the interfacial energy $E_{\rm surf}^{(i)}$ of branched segments ($i\geq 1$) and its density can be simply estimated as
$E_{\rm surf}^{(i)}=\frac{4L_{i}\gamma}{\alpha},\qquad\frac{E_{\rm surf}^{(i)}}{2h_{i}L_{i}}=\frac{2\gamma}{\alpha h_{i}},$ (7)
where $\gamma$ denotes the interfacial energy of twin boundaries (density per unit area), a material parameter that is assumed constant, i.e., independent on the interface orientation. Note that the nominal twinning planes (of normal ${\rm\bf n}$) are not necessarily perpendicular to the A–MM interface (of normal ${\rm\bf m}$), see Fig. 1, hence the factor $\alpha$ in Eq. (7), see also Eq. (5)1. Since there are two twin interfaces in each segment of the first layer ($i=0$), we have $E_{\rm surf}^{(0)}=2L_{0}\gamma/\alpha$. The actual estimate of $E_{\rm surf}^{(i)}$ used in the model of Seiner et al. (2020) includes additional correction terms that account for the deviation of the interface orientation from the theoretical one and for the different lengths of the individual interfaces within the segment. However, in our model, we will only consider the simple estimate in Eq. (7).
@@ 3.1 Twin spacing as a continuous function
In our model, the twin spacing $h$ is considered to be a continuous function of the position $x$, thus $h=h(x)$ for $0\leq x\leq L$. Considering that real microstructures do not exhibit periodicity nor perfect arrangement of branching generations, as assumed in the discrete model of Seiner et al. (2020), the twin spacing $h(x)$ should be interpreted as the average spacing measured over a representative part of the A–MM interface and determined at a given distance from it. Note that, in accord with Seiner et al. (2020), the twin spacing $h(x)$ is here measured in the direction parallel to the A–MM interface. The spatial coordinate $x$ is measured in the direction perpendicular to the A–MM interface with $x=0$ at the free boundary and $x=L$ at the A–MM interface, see Fig. 1(a).
@@ 3.2 Governing equations
The total free energy $\Phi$ involves four parameters, namely $\gamma$, $A$, $\alpha$, and $\Gamma$, and the size $L$ of the branched domain. The twin boundary energy $\gamma$ is a material parameter. Parameters $A$, $\alpha$, and $\Gamma$ depend on the specific A–MM microstructure. Moreover, $A$ and $\Gamma$ depend on the elastic properties of the phases.
@@
$0=\delta_{h}\Phi=\int_{0}^{L}\left(Ah_{,x}\delta h_{,x}-\frac{2\gamma}{\alpha h^{2}}\delta h\right){\rm{d}}x+\Gamma\,\delta h\big{|}_{x=L}\qquad\forall\,\delta h.$ (13)
@@
$h_{,xx}+\frac{2\gamma}{\alpha Ah^{2}}=0\quad\mbox{for}\;x\in[0,L]$ (15)
@@ 3.4 Model predictions
Sample results are shown in Fig. 2 in order to illustrate the predictions delivered by the model. The material parameters correspond to those adopted by Seiner et al. (2020) for the CuAlNi alloy, and the same specific A–MM microstructure is considered. Specifically, the twin boundary energy is adopted as $\gamma=0.1$ J/m2, and parameter $A$ characterizing the elastic strain energy of branching is computed according to Eq. (9)2 using $\mu=70$ GPa, $\lambda=0.6992$, $g_{2}=0.03648$, and $\alpha=0.5642$, so that $A=8.7$ MJ/m3.
@@
Finally, two ranges of the size $L$ of the twinned martensite domain are considered. The first range corresponds to a relatively large domain with $L$ varied between $1$ mm and $5$ mm, where the latter value was used in the main illustrating example of Seiner et al. (2020), see their Fig. 8. The respective results are shown in Fig. 2(a). The second range corresponds to a relatively small domain with $L$ varied between $4\,\mu$m and $20\,\mu$m, and the respective results are shown in Fig. 2(b). It can be seen that twin branching and the corresponding refinement of the microstructure towards the A–MM interface is more pronounced for larger domains. The effect of parameter $\bar{\Gamma}$ characterizing the elastic micro-strain energy at the A–MM interface is negligible in the case of relatively large domain, as shown in Fig. 2(a), except in the vicinity of the A–MM interface, however, the corresponding difference is not discernible in Fig. 2(a). In the case of the relatively small domain, Fig. 2(b), the effect of $\bar{\Gamma}$ is visible, and it is pronounced more at the A–MM interface, at $x=L$, than at the free end, at $x=0$.
@@
Fig. 4 showcases the individual contributions $\Phi_{i}$ to the total free energy $\Phi$ as a function of $\bar{\gamma}$, where
@@ 4.1 Overall dissipation potential
In the continuous setting, the twin spacing $h$ is now considered as a function of both position and time, $h=h(x,t)$. Additionally, the length of the twinned domain is time-dependent, such that $0\leq x\leq L(t)$.
We assume that, at the local level, the dissipative behavior is governed by a local rate-independent dissipation potential (density per unit interface area) that depends on the local interface propagation speed $\hat{v}_{\rm n}$ (normal to the interface),
@@ 4.3 Model predictions
For the purpose of presentation of the results, it is convenient to introduce an alternative normalization based on the maximum length $L_{\rm max}$, rather than on the current length $L$ as in Eqs. 18, 20, and (21). The corresponding dimensionless quantities are denoted by a superimposed tilde, namely
@@
Sample results are presented in Fig. 5 for $L_{\rm max}=20\,\mu$m which corresponds to $\tilde{\gamma}=10^{-3}$ for the adopted material parameters. The evolution of twin spacing $\tilde{h}$ is shown in Fig. 5(a,b) where the individual curves correspond to selected values of the domain length $\tilde{L}$. As a reference, the dashed lines show the corresponding profiles of twin spacing obtained by minimization of the total free energy (no dissipation, $\hat{r}=0$).
@@
In the second part, when the domain length $\tilde{L}$ decreases (backward motion of the A–MM interface), Fig. 5(b), the evolution is more complex and comprises two stages. In the first stage of the backward motion, the twinned domain is divided into two zones. In the outer zone, closer to the A–MM interface at $\tilde{x}=\tilde{L}$, the twin spacing evolves, while in the inner zone, closer to the free end at $\tilde{x}=0$, the twin spacing does not evolve, as the evolution is stopped by the dissipative forces. This can be seen in Fig. 5(b) where the profiles of $\tilde{h}$ corresponding to $\tilde{L}=1$ and $\tilde{L}=0.8$ overlap over a part of the domain, approximately $0\leq\tilde{x}\leq 0.4$. The no-evolution zone gradually decreases during the first stage, and when it vanishes the second stage of the backward motion starts, in which the twin spacing evolves in the entire domain.
@@
Figure 5: Evolution of twin spacing $\tilde{h}$ for $\tilde{\gamma}=10^{-3}$ ($L_{\rm max}=20\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa): (a,b) $\tilde{h}$ as a function of $\tilde{x}$ while $\tilde{L}$ is increasing (a) and decreasing (b); (c,d) twin spacing at $\tilde{x}=0$ (c) and at $\tilde{x}=\tilde{L}$ (d) as a function of the domain length $\tilde{L}$. The dashed lines correspond to the case with no dissipation, $\hat{r}=0$.
The two-stage evolution of twin spacing during the backward motion can be seen in the plots of twin spacings $\tilde{h}_{0}$ and $\tilde{h}_{L}$, belonging to the ends, see Fig. 5(c,d). The transition from the first stage of the backward motion to the second one is marked by a kink in the evolution plots. In particular, at the free end, $\tilde{x}=0$, the twin spacing $\tilde{h}_{0}$ is constant during the first stage, i.e., until $\tilde{L}\approx 0.6$ is reached, see Fig. 5(c). When $\tilde{L}$ decreases further, during the second stage, the twin spacing $\tilde{h}_{0}$ starts to decrease. At the same time, the twin spacing at the A–MM interface at $\tilde{x}=\tilde{L}$, increases during the first stage of the backward motion and decreases during the second stage. Similar kinks are also seen in the plots of the evolution of the energy contributions that are discussed next.
@@
To investigate the effect of energy dissipation further, two other cases are considered and reported in Figs. 7 and 8. Specifically, the maximum domain length $L_{\rm max}$ is decreased and increased 10 times with respect to the previous example, while all the remaining parameters remain unchanged. The case of $L_{\rm max}=2\,\mu$m, which corresponds to $\tilde{\gamma}=10^{-2}$, is shown in Fig. 7. It can be seen that the effect of dissipation on the evolution of twin spacing is more pronounced than for $\tilde{\gamma}=10^{-3}$ shown in Fig. 5. Similarly, the effect on the total free energy $\tilde{\Phi}$ is also more pronounced, see Fig. 7(c). On the other hand, for $L_{\rm max}=200\,\mu$m, which corresponds to $\tilde{\gamma}=10^{-4}$, the effect of energy dissipation on the twin spacing is visible, but small, while the effect on the total free energy is negligible, see Fig. 8.
@@
The results presented above, Figs. 58, correspond to a relatively large value of the rate-independent threshold ($\hat{r}=4$ MPa) and to relatively small domain sizes ($L_{\rm max}$ varied between 2 and 200 $\mu$m so that $\tilde{\gamma}$ varies between $10^{-2}$ and $10^{-4}$, respectively). For these parameters, the effect of energy dissipation on the evolution of branched microstructure, as characterized by the twin spacing, is visible, while the effect on the total free energy is limited (the effect is more pronounced for the individual energy contributions). For larger domain sizes $L_{\rm max}$ (i.e., for smaller values of the dimensionless twin boundary energy $\tilde{\gamma}$), the respective effects are smaller and can be considered insignificant. Of course, the effects of energy dissipation are also weaker for smaller values of the rate-independent threshold $\hat{r}$. For instance, when the case of $\tilde{\gamma}=10^{-2}$ is considered, as in Fig. 7, but parameter $\hat{r}$ is reduced twice (to 2 MPa), the effect of energy dissipation is qualitatively similar (although somewhat more pronounced) as in the case of $\tilde{\gamma}=10^{-3}$ and $\hat{r}=4$ MPa, see Figs. 5 and 6. The corresponding results are not shown for brevity.

### v4

@@ Abstract
We develop a 1D model of twin branching in shape memory alloys. The free energy of the branched microstructure comprises the interfacial and elastic strain energy contributions, both expressed in terms of the average twin spacing treated as a continuous function of the position. The total free energy is then minimized, and the corresponding Euler–Lagrange equation is solved numerically using the finite element method. The model can be considered as a continuous counterpart of the recent discrete model of Seiner et al. (2020), and our results show a very good agreement with that model in the entire range of physically relevant parameters. Furthermore, our continuous setting facilitates incorporation of energy dissipation into the model. The effect of rate-independent dissipation on the evolution of the branched microstructure is thus studied. The results show that significant effects on the microstructure and energy of the system are expected only for relatively small domain sizes. 22footnotetext: Published in Eur. J. Mech. A/Solids, 2025, doi: 10.1016/j.euromechsol.2025.105671
@@ 1 Introduction
The theoretical analysis of twin branching started with the seminal contributions of Kohn and Müller (1992, 1994), followed by many related works, mostly in the mathematics community (e.g., Conti, 2000; Capella and Otto, 2009; Chan and Conti, 2015; Dondl et al., 2016; Simon, 2021). This line of research delivers, in particular, the scaling laws for energy and for characteristic dimensions of the branched microstructure. As a follow up, a quantitative model of twin branching in shape memory alloys has been recently developed by Seiner et al. (2020). The model assumes that the branched microstructure consists of a family of self-similar branching generations, as in several earlier models, and employs refined estimates of the elastic strain energy of branching that have been derived for the considered microstructure. The parameters characterizing the actual microstructure are then obtained by minimizing the total free energy. This model will be used as a reference for our model developed in this work.
Direct spatially-resolved simulation of a branched microstructure by using, for example, the phase-field method, is not feasible because of the multiscale nature of the branched microstructure. However, some features characteristic for twin branching can be observed in phase-field simulations (e.g., Finel et al., 2010; Tůma and Stupkiewicz, 2016; Amini et al., 2023). These are mostly limited to tip splitting of martensite needles, or the like, and thus correspond to just a single branching generation. To the best of our knowledge, spatially-resolved simulations of hierarchical branched microstructures have not been reported to date. A kind of hierarchical twin branching is reported in the related work of Li and Luskin (1999), but the interfacial energy is not included in that model (only the elastic strain energy is minimized) so that the model does not capture the essential competition between the interfacial and elastic strain energy contributions.
@@
We call our model a continuous model since the twin spacing is treated as a continuous function of the position and the governing equation can be formulated as an ODE. This is in contrast to the model of Seiner et al. (2020), which we call a discrete model, since that model considers a discrete number of branching generations thus leading to a step-like dependence of the twin spacing on the position. Moreover, the solution in that model is obtained from a discrete minimization problem, in particular, involving minimization with respect to the (integer) number of branching generations.
A model similar to our model has been recently presented by Zhang et al. (2024) for the analysis of experimental data on austenite–martensite interfacial patterns in NiMnGa reported in the reference. That model also considers the twin spacing as a continuous function of the position and adopts the same form of the free energy density in the branched martensite as in our model. The difference is that our model includes the energy of elastic micro-strains at the A–MM interface, which is disregarded in the model of Zhang et al. (2024). This leads to different boundary conditions for the corresponding ODE. The model of Zhang et al. (2024), once fitted to their experimental data, correctly describes the profile of the twin spacing in the vicinity of the austenite–martensite interface.
@@
The remainder of the paper is organized as follows. In Section 2, the discrete model of twin branching of Seiner et al. (2020) is briefly described as a reference for our subsequent developments. In Section 3, the continuous model of twin branching is developed and sample results of finite-element computations are presented, including a comparison with the predictions of the model of Seiner et al. (2020). As a side note, an analytical solution to this (static) problem is discussed in A. In Section 4, energy dissipation is introduced into the model and the corresponding effects are studied for a few representative cases of evolution of the branched microstructure. Technical details concerning the derivation of the evolution problem and its computational treatment are presented in B and C, respectively. A general discussion on the model and its applicability is presented in Section 5, followed by concluding remarks in Section 6.
@@ 2.1 Austenite–twinned martensite interface
where $\hat{{\rm\bf a}}$ is a non-zero vector and ${\rm\bf R}$ is a rotation tensor. Consider further a simple laminate formed by the two martensite variants with $\lambda$ denoting the volume fraction of the variant ${\rm\bf U}_{A}$. The overall compatibility of the austenite and twin laminate is then expressed in the form of the habit plane equation,
@@ 2.2 Discrete model of twin branching
The interfacial energy $E_{\rm surf}^{(i)}$ of branched segments ($i\geq 1$) has been estimated as, see Eq. (S.9) in the supplementary material of Seiner et al. (2020),
$E_{\rm surf}^{(i)}=2\gamma(1-\lambda)h_{i}\left(\frac{2}{\varepsilon_{i}}+\lambda+\frac{1+\lambda}{2}\varepsilon_{i}\right),$ (7)
where $\gamma$ is a material parameter characterizing the interfacial energy of twin boundaries (density per unit area). In this model, the actual energy of each interface is assumed to depend on the jump of the deformation gradient at this interface, which accounts, in a simple manner, for the dependence of the interfacial energy on the interface orientation. The above expression accounts also for the changes in the area (length) of the twin interfaces within each segment as a function of the aspect ratio $2h_{i}/L_{i}$ of the segment. As discussed in Section 5, the effects related to the changes in the local interfacial energy are not significant in the considered case of a CuAlNi alloy and can be disregarded.
Assuming thus that $L_{i}\gg h_{i}$, the total interfacial energy $E_{\rm surf}^{(i)}$ of four interfaces of length $L_{i}/\alpha$ and the corresponding density can be simply estimated as
$E_{\rm surf}^{(i)}=\frac{4L_{i}\gamma}{\alpha},\qquad\frac{E_{\rm surf}^{(i)}}{2h_{i}L_{i}}=\frac{2\gamma}{\alpha h_{i}}.$ (8)
Note that the nominal twinning planes (of normal ${\rm\bf n}$) are not necessarily perpendicular to the A–MM interface (of normal ${\rm\bf m}$), see Fig. 1, hence the factor $\alpha$ in Eq. (8), see also Eq. (5)1. In fact, Eq. (8)1 can be obtained from Eq. (7) by substituting $\varepsilon_{i}$ from Eq. (4)2 and by setting $h_{i}/L_{i}=0$, which amounts to disregarding the second and third term in the parentheses in Eq. (7). Since there are two twin interfaces in each segment of the first layer ($i=0$), we have $E_{\rm surf}^{(0)}=2L_{0}\gamma/\alpha$. In our model, we will only consider the simple estimate in Eq. (8).
@@ 3.1 Twin spacing as a continuous function
In our model, the twin spacing $h$ is considered to be a continuous function of the position $x$, thus $h=h(x)$ for $0\leq x\leq L$. Considering that real microstructures do not exhibit periodicity nor perfect arrangement of branching generations, as assumed in the discrete model of Seiner et al. (2020), the twin spacing $h(x)$ should be interpreted as the average spacing measured over a representative part of the A–MM interface and determined at a given distance from it. Note that, in accord with Seiner et al. (2020), the twin spacing $h(x)$ is here measured in the direction parallel to the A–MM interface. The spatial coordinate $x$ is measured in the direction perpendicular to the A–MM interface with $x=0$ at the free boundary and $x=L$ at the A–MM interface, see Fig. 1(a). As in the discrete model, the free boundary is assumed parallel to the A–MM interface, hence the length $L$ is independent of the position along the A–MM interface. This assumption can be relaxed, as discussed in Section 5.
@@ 3.2 Governing equations
The total free energy $\Phi$ involves four parameters, namely $\gamma$, $A$, $\alpha$, and $\Gamma$, along with the size $L$ of the branched domain. The twin boundary energy $\gamma$ is a material parameter. Parameters $A$, $\alpha$, and $\Gamma$ depend on the specific A–MM microstructure. Moreover, $A$ and $\Gamma$ depend on the elastic properties of the phases.
@@
$0=\delta\Phi=\int_{0}^{L}\left(Ah_{,x}\delta h_{,x}-\frac{2\gamma}{\alpha h^{2}}\delta h\right){\rm{d}}x+\Gamma\,\delta h\big{|}_{x=L}\qquad\forall\,\delta h.$ (14)
@@
$h_{,xx}+\frac{2\gamma}{\alpha Ah^{2}}=0\quad\mbox{for}\;x\in[0,L],$ (16)
@@ 3.4 Model predictions
Sample results are shown in Fig. 2 in order to illustrate the predictions delivered by the model. The material parameters correspond to those adopted by Seiner et al. (2020) for the CuAlNi alloy, and the same type of A–MM microstructure is considered. Specifically, the twin boundary energy is adopted as $\gamma=0.1$ J/m2, and parameter $A$ characterizing the elastic strain energy of branching is computed according to Eq. (10)2 using $\mu=70$ GPa, $\lambda=0.6992$, $g_{2}=0.03648$, and $\alpha=0.5642$, so that $A=8.7$ MJ/m3.
@@
Finally, two ranges of the size $L$ of the twinned martensite domain are considered. The first range corresponds to a relatively large domain with $L$ varied between $1$ mm and $5$ mm, where the latter value was used in the main illustrating example of Seiner et al. (2020), see their Fig. 8. The respective results are shown in Fig. 2(a). The second range corresponds to a relatively small domain with $L$ varied between $4\,\mu$m and $20\,\mu$m, and the respective results are shown in Fig. 2(b). It can be seen that twin branching and the corresponding refinement of the microstructure towards the A–MM interface is more pronounced for larger domains. The effect of parameter $\bar{\Gamma}$ characterizing the elastic micro-strain energy at the A–MM interface is negligible in the case of relatively large domain, as shown in Fig. 2(a), except in the vicinity of the A–MM interface, however, the corresponding difference is not discernible in Fig. 2(a). In the case of the relatively small domain, Fig. 2(b), the effect of $\bar{\Gamma}$ is visible, and it is more pronounced at the A–MM interface, at $x=L$, than at the free end, at $x=0$.
@@
Fig. 4 showcases the individual contributions $\Phi_{k}$ to the total free energy $\Phi$ as a function of $\bar{\gamma}$, where
@@ 4.1 Overall dissipation potential
In the continuous setting, the twin spacing $h$ is thus now considered as a function of both position and time, $h=h(x,t)$. Additionally, the length of the twinned domain is time-dependent, such that $0\leq x\leq L(t)$.
Note that the length $L(t)$ is here adopted as a prescribed control parameter. Accordingly, the dissipation associated with the propagation of the A–MM interface does not intervene in the model. This dissipation would affect the evolution of $L$ as a function of some external stimuli, but this is not considered in the model. In principle, this dissipation can affect the evolution of the local microstructure of the direct A–MM interface, which, in turn, can influence the value of the energy factor $\Gamma$. This is expected to be a secondary effect (if any), and it is not considered in the model, hence $\Gamma$ is assumed to be constant.
Considering thus the propagation of twin interfaces, we assume that, at the local level, the dissipative behavior is governed by a local rate-independent dissipation potential (density per unit interface area) that depends on the local interface propagation speed $\hat{v}_{\rm n}$ (normal to the interface),
@@ 4.3 Model predictions
For the purpose of presentation of the results, it is convenient to introduce an alternative normalization based on the maximum length $L_{\rm max}$, rather than on the current length $L$ as in Eqs. 19, 21, and (22). The corresponding dimensionless quantities are denoted by a superimposed tilde, namely
@@
Sample results are presented in Fig. 5 for $L_{\rm max}=20\,\mu$m which corresponds to $\tilde{\gamma}=10^{-3}$ for the adopted material parameters. The evolution of twin spacing $\tilde{h}$ is shown in Fig. 5(a,b) where the individual curves correspond to selected values of the domain length $\tilde{L}$. As a reference, the dashed lines show the profiles of twin spacing obtained by minimization of the total free energy (no dissipation, $\hat{r}=0$) for $\tilde{L}=0.04$ and for $\tilde{L}=1$ (the difference is significant only for $\tilde{L}=1$).
@@
In the second part, when the domain length $\tilde{L}$ decreases (backward motion of the A–MM interface), Fig. 5(b), the evolution is more complex and comprises two stages. In the first stage of the backward motion, the twinned domain is divided into two zones. In the outer zone, closer to the A–MM interface at $\tilde{x}=\tilde{L}$, the twin spacing evolves, while in the inner zone, closer to the free end at $\tilde{x}=0$, the twin spacing does not evolve, as the evolution is stopped by the dissipative forces. This can be seen in Fig. 5(b) where the subsequent profiles overlap over a part of the domain. The no-evolution zone gradually decreases during the first stage, and when it vanishes the second stage of the backward motion starts, in which the twin spacing evolves in the entire domain.
@@
Figure 5: Evolution of twin spacing $\tilde{h}$ for $\tilde{\gamma}=10^{-3}$ ($L_{\rm max}=20\,\mu$m) and $\bar{r}=0.21$ ($\hat{r}=4$ MPa): (a,b) $\tilde{h}$ as a function of $\tilde{x}$ while $\tilde{L}$ is increasing (a) and decreasing (b); (c,d) twin spacing at $\tilde{x}=0$ (c) and at $\tilde{x}=\tilde{L}$ (d) as a function of the domain length $\tilde{L}$. The dashed lines correspond to the case with no dissipation, $\hat{r}=0$ (in panels (a) and (b), the reference dashed lines are provided only for the smallest and largest $\tilde{L}$).
The two-stage evolution of twin spacing during the backward motion can be seen in the plots of twin spacings $\tilde{h}_{0}$ and $\tilde{h}_{L}$, belonging to the ends, see Fig. 5(c,d). The transition from the first stage of the backward motion to the second one is continuous but non-smooth, hence it is marked by a kink in the evolution plots (this is a consequence of adopting the non-smooth dissipation potential (27), and the kink corresponds to an abrupt change of the rate of evolution). In particular, at the free end, $\tilde{x}=0$, the twin spacing $\tilde{h}_{0}$ is constant during the first stage, i.e., until $\tilde{L}\approx 0.6$ is reached, see Fig. 5(c). When $\tilde{L}$ decreases further, during the second stage, the twin spacing $\tilde{h}_{0}$ starts to decrease. At the same time, the twin spacing at the A–MM interface at $\tilde{x}=\tilde{L}$, increases during the first stage of the backward motion and decreases during the second stage. Similar kinks are also seen in the plots of the evolution of the energy contributions that are discussed next.
@@
To investigate the effect of energy dissipation further, two other cases are considered and reported in Figs. 7 and 8. Specifically, the maximum domain length $L_{\rm max}$ is decreased and increased 10 times with respect to the previous example, while all the remaining parameters remain unchanged. The case of $L_{\rm max}=2\,\mu$m, which corresponds to $\tilde{\gamma}=10^{-2}$, is shown in Fig. 7. It can be seen that the effect of dissipation on the evolution of twin spacing is in this case more pronounced than for $\tilde{\gamma}=10^{-3}$ shown in Fig. 5. Similarly, the effect on the total free energy $\tilde{\Phi}$ is also more pronounced, see Fig. 7(c). On the other hand, for $L_{\rm max}=200\,\mu$m, which corresponds to $\tilde{\gamma}=10^{-4}$, the effect of energy dissipation on the twin spacing is visible, but small, while the effect on the total free energy is negligible, see Fig. 8.
@@
The results presented in Figs. 58, correspond to a relatively large value of the rate-independent threshold ($\hat{r}=4$ MPa) and to relatively small domain sizes ($L_{\rm max}$ varied between 2 and 200 $\mu$m so that $\tilde{\gamma}$ varies between $10^{-2}$ and $10^{-4}$, respectively). For these parameters, the effect of energy dissipation on the evolution of branched microstructure, as characterized by the twin spacing, is visible, while the effect on the total free energy is limited (the effect is more pronounced for the individual energy contributions). For larger domain sizes $L_{\rm max}$ (i.e., for smaller values of the dimensionless twin boundary energy $\tilde{\gamma}$), the respective effects are smaller and can be considered insignificant. Of course, the effects of energy dissipation are also weaker for smaller values of the rate-independent threshold $\hat{r}$. For instance, when the case of $\tilde{\gamma}=10^{-2}$ is considered, as in Fig. 7, but parameter $\hat{r}$ is reduced twice (to 2 MPa), the effect of energy dissipation is qualitatively similar (although somewhat more pronounced) to the case of $\tilde{\gamma}=10^{-3}$ and $\hat{r}=4$ MPa, see Figs. 5 and 6. The corresponding results are not shown.